OpenStudy (anonymous):

4x^2 =12 solve using radicals,rationalize all denominators express complex numbers in terms of i or use a comma to separate answers if needed then find the x intercepts

OpenStudy (zarkon):

divide by 4 from both sides...take square root of both sides

OpenStudy (anonymous):

ok well is this for the first part or the second part and can you show me what you mean please

OpenStudy (zarkon):

\[4x^2=12\] \[\frac{4x^2}{4}=\frac{12}{4}\] \[x^2=3\] \[x=\pm\sqrt{3}\]

OpenStudy (anonymous):

so there are two solutions to thisa problem ?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

ok whar the other one\[\sqrt{-3}\]

OpenStudy (zarkon):

\[\sqrt{3}\cdot i\]

OpenStudy (zarkon):

?

OpenStudy (anonymous):

i see sorry i was going by some past answers im really struggling with this

OpenStudy (anonymous):

wait i tried your answer and then it said mine was right but i write it wrong\[-\sqrt{3}\]

OpenStudy (zarkon):

there are two answers to the original problem you posted \[\sqrt{3}\] and \[-\sqrt{3}\] as i had posted above \[\pm\sqrt{3}\]

OpenStudy (anonymous):

i misunderstood im sorry

OpenStudy (zarkon):

That's ok :)

OpenStudy (anonymous):

so if i were to apply what you showed me then 5x^2=10 would be similar

OpenStudy (zarkon):

yes..divide by 5 first then take square root you will get two solutions

OpenStudy (anonymous):

i would get \[\sqrt{2},-\sqrt{2}\]

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

then my intercepts should be \[\sqrt{2},0,-\sqrt{2},0\]

OpenStudy (zarkon):

yes...if you wrote your equation as \[y=5x^2-10\] then the x-intercepts would be \[(\sqrt{2},0),(-\sqrt{2},0)\]

OpenStudy (anonymous):

so i need the parenthesis for them thats why i was getting it wrong?

OpenStudy (zarkon):

prob they are points

OpenStudy (anonymous):

yep that was the issue lol

OpenStudy (zarkon):

ah